# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 2.2. $\infty$-operads and $\infty$-cooperads
Let ${\mathcal{C}}$ be a preadditive symmetric monoidal $\infty$-category compatible with small colimits that admits small limits. For any non-counital $\infty$-cooperad ${\mathcal{Q}}\in\mathrm{CoOp}({\mathcal{C}})^{\mathrm{nu}}\subset{\mathrm{coAlg}}({\mathrm{sSeq}}({\mathcal{C}}))$ we set
${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{\mathcal{Q}}({\mathcal{C}}):={\mathrm{coLMod}}_{\mathcal{Q}}({\mathcal{C}}),$
@@
Let ${\mathcal{C}}$ be a preadditive symmetric monoidal $\infty$-category compatible with small colimits that admits small limits. For any non-counital $\infty$-cooperad ${\mathcal{Q}}\in\mathrm{CoOp}({\mathcal{C}})^{\mathrm{nu}}\subset{\mathrm{coAlg}}({\mathrm{sSeq}}({\mathcal{C}}))$ there is a forgetful functor over ${\mathcal{C}}:$
${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}}).$
@@
${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})=\mathrm{coLMod}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathrm{LMod}}_{{\mathcal{Q}}}({\mathcal{C}}^{\mathrm{op}})^{\mathrm{op}}={\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}}),$
@@
    Let ${\mathcal{C}}$ be a preadditive symmetric monoidal $\infty$-category compatible with small colimits that admits small limits and ${\mathcal{Q}}$ a non-counital $\infty$-cooperad in ${\mathcal{C}}.$ The forgetful functor ${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathcal{C}}$ is comonadic.
@@
1.: The $\infty$-category $\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})$ is the $\infty$-category of left modules with respect to a left action of a monoidal $\infty$-category and so is monadic [23, Example 4.7.3.9.].
2.: We embed ${\mathcal{C}}$ symmetric monoidally into ${\mathcal{P}}({\mathcal{C}})$ endowed with Day-convolution, which is compatible with small colimits. By 1. the forgetful functor $\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{P}}({\mathcal{C}}))\to{\mathcal{P}}({\mathcal{C}})$ is monadic. The forgetful functor $\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\to{\mathcal{C}}$ is the pullback of the latter monadic forgetful functor. This implies that every simplicial object in $\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})$ whose image in ${\mathcal{C}}$ splits, admits a geometric realization that is preserved by the forgetful functor. Hence the claim follows from the Barr-Beck theorem [23, Theorem 4.7.3.5].
3.: The $\infty$-category ${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})$ is the $\infty$-category of left comodules with respect to a left action of a monoidal $\infty$-category. Dualising [23, Example 4.7.3.9.] the $\infty$-category of left comodules with respect to a left action of a monoidal $\infty$-category is comonadic.
@@
Let ${\mathcal{C}}$ be a preadditive symmetric monoidal $\infty$-category compatible with small colimits and ${\mathcal{Q}}$ a non-counital $\infty$-cooperad in ${\mathcal{C}}.$ The $\infty$-category ${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})$ admits small colimits that are preserved by the forgetful functor to ${\mathcal{C}}$ by [23, Corollary 4.2.3.3.] because ${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})$ is the $\infty$-category of left comodules with respect to a left action of a monoidal $\infty$-category.
Let ${\mathcal{C}}$ be a preadditive presentably symmetric monoidal $\infty$-category. The $\infty$-category ${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})$ is presentable [20, Proposition 5.51.] since the functor ${\mathrm{coAlg}}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathcal{C}}$ is comonadic by Lemma 2.30 (3) and the associated comonad ${\mathcal{Q}}\circ(-):{\mathcal{C}}\to{\mathcal{C}}$ is accessible.
@@
Let ${\mathcal{C}}$ be a symmetric monoidal $\infty$-category having small colimits and ${\mathrm{K}}\to{\mathrm{Op}}({\mathcal{C}})$ a functor such that the composition ${\mathrm{K}}\to{\mathrm{Op}}({\mathcal{C}})\to{\mathrm{sSeq}}({\mathcal{C}})$ admits a limit. Then the functor ${\mathrm{K}}\to{\mathrm{Op}}({\mathcal{C}})$ has a limit, which is preserved by the forgetful functor ${\mathrm{Op}}({\mathcal{C}})\to{\mathrm{sSeq}}({\mathcal{C}}).$ And similarly, for ${\mathrm{Op}}({\mathcal{C}})$ replaced by ${\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}$.
@@
We can generally assume that ${\mathrm{K}}$ is small by changing to a larger universe if necessary. By Proposition 7.9 there is a symmetric monoidal embedding ${\mathcal{C}}\subset{\mathcal{C}}^{\prime}$ into a symmetric monoidal $\infty$-category compatible with small colimits that admits small limits such that the embedding preserves small limits. The embedding ${\mathcal{C}}\subset{\mathcal{C}}^{\prime}$ induces embeddings ${\mathrm{Op}}({\mathcal{C}})\subset{\mathrm{Op}}({\mathcal{C}}^{\prime}),{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}\subset{\mathrm{Op}}({\mathcal{C}}^{\prime})^{\mathrm{nu}}$. Consequently, we can assume that the symmetric monoidal structure on ${\mathcal{C}}$ is compatible with small colimits. In this case the result follows from [23, Corollary 3.2.2.4.] since ${\mathrm{Op}}({\mathcal{C}})\simeq\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}})),{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}\simeq\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})$ are $\infty$-categories of associative algebras in a monoidal $\infty$-category.
@@ 2.3. Shifting $\infty$-operads
In this subsection we construct a shift functor on the $\infty$-category of $\infty$-operads and $\infty$-cooperads in any stable symmetric monoidal $\infty$-category ${\mathcal{C}}$ having small colimits. In the following let []delimited-[][-][ - ] be the (object-wise) shift functor of ${\mathrm{sSeq}}({\mathcal{C}}).$
@@
Let ${\mathcal{C}}$ be a stable symmetric monoidal $\infty$-category compatible with small colimits and ${\mathrm{n}}\in{\mathbb{Z}}.$ Then ${\mathrm{triv}}[{\mathrm{n}}]$ is tensor inverse to ${\mathrm{triv}}[-{\mathrm{n}}]$ in the composition product.
@@
For every ${\mathrm{n}},{\mathrm{m}}\in{\mathbb{Z}}$ and ${\mathrm{X}},{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})$ such that ${\mathrm{X}}$ is concentrated in degree 1 there is a canonical equivalence
${\mathrm{X}}[{\mathrm{n}}]\circ{\mathrm{Y}}[{\mathrm{m}}]\simeq\coprod_{{\mathrm{k}}\in{\mathbb{N}}}({\mathrm{X}}[{\mathrm{n}}])_{\mathrm{k}}\otimes_{\Sigma_{\mathrm{k}}}({\mathrm{Y}}[{\mathrm{m}}])^{\otimes{\mathrm{k}}}\simeq{\mathrm{X}}_{1}[{\mathrm{n}}]\otimes{\mathrm{Y}}[{\mathrm{m}}]\simeq({\mathrm{X}}_{1}\otimes{\mathrm{Y}})[{\mathrm{n}}+{\mathrm{m}}].$
In particular, there is a canonical equivalence ${\mathrm{triv}}[{\mathrm{n}}]\circ{\mathrm{triv}}[{\mathrm{m}}]\simeq{\mathrm{triv}}[{\mathrm{n}}+{\mathrm{m}}].$
@@
Let ${\mathcal{C}}$ be a stable symmetric monoidal $\infty$-category compatible with small colimits and ${\mathrm{n}}\in{\mathbb{Z}}.$ Via the canonical equivalence $\mathrm{Fun}^{\otimes,{\mathrm{L}}}_{{\mathcal{C}}/}({\mathrm{sSeq}}({\mathcal{C}}),{\mathrm{sSeq}}({\mathcal{C}}))\simeq{\mathrm{sSeq}}({\mathcal{C}})$ evaluating at ${\mathrm{triv}}$ the tensor invertible object ${\mathrm{triv}}[{\mathrm{n}}]$ corresponds to a symmetric monoidal autoequivalence $\alpha_{\mathrm{n}}$ of ${\mathrm{sSeq}}({\mathcal{C}})$ under ${\mathcal{C}}$, which gives rise to a monoidal autoequivalence of $\mathrm{Fun}^{\otimes,{\mathrm{L}}}_{{\mathcal{C}}/}({\mathrm{sSeq}}({\mathcal{C}}),{\mathrm{sSeq}}({\mathcal{C}}))\simeq{\mathrm{sSeq}}({\mathcal{C}})$ conjugating with $\alpha_{\mathrm{n}}$ denoted by $(-)({\mathrm{n}})$.
@@
Let ${\mathcal{C}}$ be a stable symmetric monoidal $\infty$-category compatible with small colimits. For every ${\mathrm{n}}\in{\mathbb{Z}},{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})$ and ${\mathrm{r}}\in\Sigma$ there is a natural equivalence
${\mathrm{Y}}({\mathrm{n}})_{\mathrm{r}}\simeq{\mathrm{Y}}_{\mathrm{r}}[(1-{\mathrm{r}}){\mathrm{n}}].$
@@
${\mathrm{Y}}({\mathrm{n}})\simeq(\alpha_{-{\mathrm{n}}}\circ\bar{{\mathrm{Y}}}\circ\alpha_{\mathrm{n}})({\mathrm{triv}})\simeq\alpha_{-{\mathrm{n}}}(\bar{{\mathrm{Y}}}({\mathrm{triv}}[{\mathrm{n}}]))\simeq\alpha_{-{\mathrm{n}}}(\bar{{\mathrm{Y}}}({\mathrm{triv}}))[{\mathrm{n}}]$
$\simeq\alpha_{-{\mathrm{n}}}({\mathrm{Y}})[{\mathrm{n}}]\simeq({\mathrm{Y}}\circ({\mathrm{triv}}[-{\mathrm{n}}]))[{\mathrm{n}}]\simeq$
$(\coprod_{{\mathrm{k}}\in\Sigma}{\mathrm{Y}}_{\mathrm{k}}\otimes_{\Sigma_{\mathrm{k}}}{\mathrm{triv}}[-{\mathrm{n}}]^{\otimes{\mathrm{k}}})[{\mathrm{n}}]\simeq\coprod_{{\mathrm{k}}\in\Sigma}{\mathrm{Y}}_{\mathrm{k}}\otimes_{\Sigma_{\mathrm{k}}}{\mathrm{triv}}^{\otimes{\mathrm{k}}}[(1-{\mathrm{k}}){\mathrm{n}}]$
and thus ${\mathrm{Y}}({\mathrm{n}})_{\mathrm{r}}\simeq{\mathrm{Y}}_{\mathrm{r}}\otimes_{\Sigma_{\mathrm{r}}}(\Sigma_{\mathrm{r}}\times{\mathbb{1}})[(1-{\mathrm{r}}){\mathrm{n}}]\simeq{\mathrm{Y}}_{\mathrm{r}}[(1-{\mathrm{r}}){\mathrm{n}}].$
@@
Lemma 2.35 shows that for any ${\mathrm{n}}\in{\mathbb{Z}}$ the autoequivalence $(-)({\mathrm{n}})$ of ${\mathrm{sSeq}}({\mathcal{C}})$ restricts to the ${\mathrm{n}}$-fold shift of ${\mathcal{C}}$ along the embedding ${\mathcal{C}}\subset{\mathrm{sSeq}}({\mathcal{C}})$ viewing an object concentrated in degree zero. Moreover $(-)({\mathrm{n}})$ restricts to an autoequivalence of ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}.$
The monoidal autoequivalence $(-)({\mathrm{n}})$ of ${\mathrm{sSeq}}({\mathcal{C}})$ gives rise to induced autoequivalences of ${\mathrm{coAlg}}({\mathrm{sSeq}}({\mathcal{C}})),{\mathrm{coLMod}}({\mathcal{C}}).$ If ${\mathcal{C}}$ is preadditive, for every non-counital $\infty$-cooperad ${\mathcal{Q}}$ in ${\mathcal{C}}$ we obtain a commutative square
[…]
@@
If ${\mathcal{C}}$ is a stable symmetric monoidal $\infty$-category having small colimits, Corollary 7.10 states that ${\mathcal{C}}$ embeds symmetric monoidally and exact into a stable symmetric monoidal $\infty$-category ${\mathcal{C}}^{\prime}$ compatible with small colimits. Thus the non-symmetric $\infty$-operad ${\mathrm{sSeq}}({\mathcal{C}})^{\otimes}$ embeds into the monoidal $\infty$-category ${\mathrm{sSeq}}({\mathcal{C}}^{\prime})^{\otimes}$ and for any ${\mathrm{n}}\in{\mathbb{Z}}$ the monoidal autoequivalence $(-)({\mathrm{n}})$ of ${\mathrm{sSeq}}({\mathcal{C}}^{\prime})$ restricts to a monoidal autoequivalence $(-)({\mathrm{n}})$ of ${\mathrm{sSeq}}({\mathcal{C}})$. The autoequivalence $(-)({\mathrm{n}})$ gives rise to autoequivalences of ${\mathrm{Op}}({\mathcal{C}})=\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}}))$ and ${\mathrm{LMod}}({\mathcal{C}}).$ For every $\infty$-operad ${\mathcal{O}}$ in ${\mathcal{C}}$ we obtain a commutative square
[…]

### v2

@@ 2.1. The composition product
Let $\mathcal{C}$} be a preadditive symmetric monoidal $\infty$-category compatible with small colimits that admits small limits. For any non-counital $\infty$-cooperad ${\mathcal{Q}}\in\mathrm{CoOp}({\mathcal{C}})^{\mathrm{nun}}\subset\mathrm{co}\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}}))$ we set
$\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{\mathcal{Q}}({\mathcal{C}}):=\mathrm{co}{\mathrm{LMod}}_{\mathcal{Q}}({\mathcal{C}}),$
@@
Let $\mathcal{C}$} be a preadditive symmetric monoidal $\infty$-category compatible with small colimits that admits small limits. For any non-counital $\infty$-cooperad ${\mathcal{Q}}\in\mathrm{CoOp}({\mathcal{C}})^{\mathrm{nun}}\subset\mathrm{co}\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}}))$ there is a forgetful functor over $\mathcal{C}$}:
$\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}}).$
@@
$\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})=\mathrm{coLMod}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathrm{LMod}}_{{\mathcal{Q}}}({\mathcal{C}}^{\mathrm{op}})^{\mathrm{op}}=\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}}),$
@@
    Let $\mathcal{C}$} be a preadditive symmetric monoidal $\infty$-category compatible with small colimits that admits small limits and $\mathcal{Q}$} a non-counital $\infty$-cooperad in $\mathcal{C}$}. The forgetful functor $\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathcal{C}}$ is comonadic.
@@
1.: The $\infty$-category $\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})$ is the $\infty$-category of left modules with respect to a left action of a monoidal $\infty$-category and so is monadic [39, Example 4.7.3.9.].
2.: We embed $\mathcal{C}$} symmetric monoidally into ${\mathcal{P}}({\mathcal{C}})$ endowed with Day-convolution, which is compatible with small colimits. By 1. the forgetful functor $\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{P}}({\mathcal{C}}))\to{\mathcal{P}}({\mathcal{C}})$ is monadic. The forgetful functor $\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\to{\mathcal{C}}$ is the pullback of the latter monadic forgetful functor. This implies that every simplicial object in $\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})$ whose image in $\mathcal{C}$} splits, admits a geometric realization that is preserved by the forgetful functor. Hence the claim follows from the Barr-Beck theorem [39, Theorem 4.7.3.5].
3.: The $\infty$-category $\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})$ is the $\infty$-category of left comodules with respect to a left action of a monoidal $\infty$-category. Dualising [39, Example 4.7.3.9.] the $\infty$-category of left comodules with respect to a left action of a monoidal $\infty$-category is comonadic.
@@
Let $\mathcal{C}$} be a preadditive symmetric monoidal $\infty$-category compatible with small colimits and $\mathcal{Q}$} a non-counital $\infty$-cooperad in $\mathcal{C}$}. The $\infty$-category $\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})$ admits small colimits that are preserved by the forgetful functor to $\mathcal{C}$} by [39, Corollary 4.2.3.3.] because $\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})$ is the $\infty$-category of left comodules with respect to a left action of a monoidal $\infty$-category.
Let $\mathcal{C}$} be a preadditive presentably symmetric monoidal $\infty$-category. The $\infty$-category $\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})$ is presentable [23, Proposition 5.51.] since the functor $\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}}({\mathcal{C}})\to{\mathcal{C}}$ is comonadic by Lemma 2.30 (3) and the associated comonad ${\mathcal{Q}}\circ(-):{\mathcal{C}}\to{\mathcal{C}}$ is accessible.
@@
Let $\mathcal{C}$} be a symmetric monoidal $\infty$-category having small colimits and ${\mathrm{K}}\to{\mathrm{Op}}({\mathcal{C}})$ a functor such that the composition ${\mathrm{K}}\to{\mathrm{Op}}({\mathcal{C}})\to{\mathrm{sSeq}}({\mathcal{C}})$ admits a limit. Then the functor ${\mathrm{K}}\to{\mathrm{Op}}({\mathcal{C}})$ has a limit, which is preserved by the forgetful functor ${\mathrm{Op}}({\mathcal{C}})\to{\mathrm{sSeq}}({\mathcal{C}}).$ And similarly, for ${\mathrm{Op}}({\mathcal{C}})$ replaced by ${\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}$.
@@
We can generally assume that $\mathrm{K}$} is small by changing to a larger universe if necessary. By Proposition 5.9 there is a symmetric monoidal embedding ${\mathcal{C}}\subset{\mathcal{C}}^{\prime}$ into a symmetric monoidal $\infty$-category compatible with small colimits that admits small limits such that the embedding preserves small limits. The embedding ${\mathcal{C}}\subset{\mathcal{C}}^{\prime}$ induces embeddings ${\mathrm{Op}}({\mathcal{C}})\subset{\mathrm{Op}}({\mathcal{C}}^{\prime}),{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}\subset{\mathrm{Op}}({\mathcal{C}}^{\prime})^{\mathrm{nun}}$. Consequently, we can assume that the symmetric monoidal structure on $\mathcal{C}$} is compatible with small colimits. In this case the result follows from [39, Corollary 3.2.2.4.] since ${\mathrm{Op}}({\mathcal{C}})\simeq\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}})),{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}\simeq\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})$ are $\infty$-categories of associative algebras in a monoidal $\infty$-category.
@@ 2.3. Shifting $\infty$-operads
In this subsection we construct a shift functor on the $\infty$-category of $\infty$-operads and $\infty$-cooperads in any stable symmetric monoidal $\infty$-category $\mathcal{C}$} having small colimits. In the following let [][-] be the (object-wise) shift functor of ${\mathrm{sSeq}}({\mathcal{C}}).$
@@
Let $\mathcal{C}$} be a stable symmetric monoidal $\infty$-category compatible with small colimits and ${n}\in{\mathbb{Z}}.$ Then ${\mathrm{triv}}[{n}]$ is tensor inverse to ${\mathrm{triv}}[-{n}]$ in the composition product.
@@
For every ${n},{m}\in{\mathbb{Z}}$ and ${\mathrm{X}},{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})$ such that $\mathrm{X}$} is concentrated in degree 1 there is a canonical equivalence
${\mathrm{X}}[{n}]\circ{\mathrm{Y}}[{m}]\simeq\coprod_{{k}\in{\mathbb{N}}}({\mathrm{X}}[{n}])_{k}\otimes_{\Sigma_{k}}({\mathrm{Y}}[{m}])^{\otimes{k}}\simeq{\mathrm{X}}_{1}[{n}]\otimes{\mathrm{Y}}[{m}]\simeq({\mathrm{X}}_{1}\otimes{\mathrm{Y}})[{n}+{m}].$
In particular, there is a canonical equivalence ${\mathrm{triv}}[{n}]\circ{\mathrm{triv}}[{m}]\simeq{\mathrm{triv}}[{n}+{m}].$
@@
Let $\mathcal{C}$} be a stable symmetric monoidal $\infty$-category compatible with small colimits and ${n}\in{\mathbb{Z}}.$ Via the canonical equivalence $\mathrm{Fun}^{\otimes,{\mathrm{L}}}_{{\mathcal{C}}/}({\mathrm{sSeq}}({\mathcal{C}}),{\mathrm{sSeq}}({\mathcal{C}}))\simeq{\mathrm{sSeq}}({\mathcal{C}})$ evaluating at $\mathrm{triv}$} the tensor invertible object ${\mathrm{triv}}[{n}]$ corresponds to a symmetric monoidal autoequivalence $\alpha_{n}$ of ${\mathrm{sSeq}}({\mathcal{C}})$ under $\mathcal{C}$}, which gives rise to a monoidal autoequivalence of $\mathrm{Fun}^{\otimes,{\mathrm{L}}}_{{\mathcal{C}}/}({\mathrm{sSeq}}({\mathcal{C}}),{\mathrm{sSeq}}({\mathcal{C}}))\simeq{\mathrm{sSeq}}({\mathcal{C}})$ conjugating with $\alpha_{n}$ denoted by ()(n)(-)({n}).
@@
Let $\mathcal{C}$} be a stable symmetric monoidal $\infty$-category compatible with small colimits. For every ${n}\in{\mathbb{Z}},{\mathrm{Y}}\in{\mathrm{sSeq}}({\mathcal{C}})$ and ${r}\in\Sigma$ there is a natural equivalence
${\mathrm{Y}}({n})_{r}\simeq{\mathrm{Y}}_{r}[(1-{r}){n}].$
@@
${\mathrm{Y}}({n})\simeq(\alpha_{-{n}}\circ\bar{{\mathrm{Y}}}\circ\alpha_{n})({\mathrm{triv}})\simeq\alpha_{-{n}}(\bar{{\mathrm{Y}}}({\mathrm{triv}}[{n}]))\simeq\alpha_{-{n}}(\bar{{\mathrm{Y}}}({\mathrm{triv}}))[{n}]$
$\simeq\alpha_{-{n}}({\mathrm{Y}})[{n}]\simeq({\mathrm{Y}}\circ({\mathrm{triv}}[-{n}]))[{n}]\simeq$
$(\coprod_{{k}\in\Sigma}{\mathrm{Y}}_{k}\otimes_{\Sigma_{k}}{\mathrm{triv}}[-{n}]^{\otimes{k}})[{n}]\simeq\coprod_{{k}\in\Sigma}{\mathrm{Y}}_{k}\otimes_{\Sigma_{k}}{\mathrm{triv}}^{\otimes{k}}[(1-{k}){n}]$
and thus ${\mathrm{Y}}({n})_{r}\simeq{\mathrm{Y}}_{r}\otimes_{\Sigma_{r}}(\Sigma_{r}\times{\mathbb{1}})[(1-{r}){n}]\simeq{\mathrm{Y}}_{r}[(1-{r}){n}].$
@@
Lemma 2.35 shows that for any ${n}\in{\mathbb{Z}}$ the autoequivalence ()(n)(-)({n}) of ${\mathrm{sSeq}}({\mathcal{C}})$ restricts to the n{n}-fold shift of $\mathcal{C}$} along the embedding ${\mathcal{C}}\subset{\mathrm{sSeq}}({\mathcal{C}})$ viewing an object concentrated in degree zero. Moreover ()(n)(-)({n}) restricts to an autoequivalence of ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}.$
The monoidal autoequivalence ()(n)(-)({n}) of ${\mathrm{sSeq}}({\mathcal{C}})$ gives rise to induced autoequivalences of $\mathrm{co}\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}})),\mathrm{co}{\mathrm{LMod}}({\mathcal{C}}).$ If $\mathcal{C}$} is preadditive, for every non-counital $\infty$-cooperad $\mathcal{Q}$} in $\mathcal{C}$} we obtain a commutative square
${\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{{\mathcal{Q}}({n})}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\mathrm{co}\mathrm{Alg}^{\mathrm{dp},{\mathrm{conil}}}_{\mathcal{Q}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{C}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(-)[-{n}]}\scriptstyle{\simeq}$${{\mathcal{C}}.}$
@@
If $\mathcal{C}$} is a stable symmetric monoidal $\infty$-category having small colimits, Corollary 5.11 states that $\mathcal{C}$} embeds symmetric monoidally and exact into a stable symmetric monoidal $\infty$-category $\mathcal{C}$$^{\prime}$ compatible with small colimits. Thus the non-symmetric $\infty$-operad ${\mathrm{sSeq}}({\mathcal{C}})^{\otimes}$ embeds into the monoidal $\infty$-category ${\mathrm{sSeq}}({\mathcal{C}}^{\prime})^{\otimes}$ and for any ${n}\in{\mathbb{Z}}$ the monoidal autoequivalence ()(n)(-)({n}) of ${\mathrm{sSeq}}({\mathcal{C}}^{\prime})$ restricts to a monoidal autoequivalence ()(n)(-)({n}) of ${\mathrm{sSeq}}({\mathcal{C}})$. The autoequivalence ()(n)(-)({n}) gives rise to autoequivalences of ${\mathrm{Op}}({\mathcal{C}})=\mathrm{Alg}({\mathrm{sSeq}}({\mathcal{C}}))$ and ${\mathrm{LMod}}({\mathcal{C}}).$ For every $\infty$-operad $\mathcal{O}$} in $\mathcal{C}$} we obtain a commutative square
${\mathrm{Alg}_{{\mathcal{O}}({n})}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{C}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(-)[-{n}]}\scriptstyle{\simeq}$${{\mathcal{C}}.}$